The limiting angle of resolution for the eye is 1.4 x 10^-4 rad. What is the minimum separation between two point sources of light that the eye can distinguish if they are at a distance of 40m from the observer?
Answer #1

minimum separation = 1.4*10-4*40 = 5.6*10-3 metre = 5.6 mili metre